# JUHAN ARU, PHILÉMON BORDEREAU

### v1

@@ Abstract.
One way to uniquely define Schramm-Loewner Evolution (SLE) in multiply connected domains is to use the restriction property. This gives an implicit definition of a $\sigma$-finite measure on curves; yet it is in general not clear how to construct such measures nor whether the mass of these measures, called the partition function, is finite.
We provide an explicit construction of the such conformal restriction SLEs in multiply connected domains when $\kappa=4$ using the Gaussian Free Field (GFF). In particular, both when the target points of the curve are on the same or on distinct boundary components, we show that there is a mixture of laws of level lines of GFFs that satisfies the restriction property. This allows us to give an expression for the partition function of $\mathrm{SLE}_{4}$ on multiply connected domains and shows that the partition function is finite, answering the question raised in [Law09, Law11].
In a second part, we provide a second construction of $\mathrm{SLE}_{\kappa}$ in multiply-connected domains for the whole range $\kappa\in(8/3,4]$; specific, however, to the case of the two target points belonging to the same boundary components. This is inspired by [WW13] and consists of a mixture of laws on curves obtained by following $\operatorname{CLE}_{\kappa}$ loops and restriction hulls attached to parts of the boundary of the domain. In this case as well, we obtain as a corollary the finiteness of the partition function for this type of $\mathrm{SLE}_{\kappa}$.
@@ Introduction
Schramm-Loewner Evolution $\mathrm{SLE}_{\kappa}$, $\kappa\geq 0$ is a one-dimensional family of conformally invariant probability measures on non-crossing curves in domains of the complex plane [Sch00] that are either proved or conjectured to describe the scaling limit of interfaces of a variety of planar statistical physics models, e.g. [Smi10, LSW04, SS09, DCLQ21, CDCH+14, KW10, She06]. Another active area of research since the introduction of SLE has been to identify and describe its connections with Conformal Field Theory (CFT), where $\mathrm{SLE}_{\kappa}$ is expected to be connected with CFTs of central charge $\mathbf{c}=\mathbf{c}(\kappa)=(6-\kappa)(3\kappa-8)/(2\kappa)$, see e.g. [Pel19] for a review.
Schramm constructed (chordal) $\mathrm{SLE}_{\kappa}$ in simply connected domains as a solution to Loewner’s equation driven by a multiple of Brownian motion $\sqrt{\kappa}B$ and showed via a short argument that the resulting law is the unique law on simple curves (modulo increasing time-reparametrization) that satisfies conformal invariance and a domain Markov property [Sch00]. Neither this construction nor the characterisation generalise directly to multiply connected domains as the conformal class changes with the growth of the curve and therefore the remaining slit domain cannot be related to the initial domain via conformal transformations–in other words, due to the non-triviality of the moduli space for more general domains, in order to identify a family of laws on curves indexed by the moduli space, one needs to specify how these laws between different conformal classes compare.
Still, variants of $\mathrm{SLE}_{\kappa}$ have been constructed in doubly connected domains via the annulus Loewner equation using the modulus of the slit domain as the time parametrization [Zha04, Zha08, HBB10, HLD08], and in higher connectivity using further generalizations of the Loewner equation [BF08, BF04]. In these latter works one also needs to keep track of the evolution of the moduli as the curve grows to obtain a Markov property and this has a downside that the natural conditions do not allow to single out an unique law on the curves, but rather one identifies a whole family of possible curves as $\mathrm{SLE}_{\kappa}$. A different approach that does single out an unique law of $\mathrm{SLE}_{\kappa}$ was suggested by Lawler [Law09] and is described next.
@@
In [LSW03], it was noted that for $\kappa\leq 4$ (the values of $\kappa$ for which the resulting curve is a.s. simple [RS05]), the law $\nu_{\mathbb{H}\setminus K,\kappa}(0,+\infty)$ of $\mathrm{SLE}_{\kappa}$ from 00 to $\infty$ with $K$ compact, $\mathrm{dist}(K,0)>0$ and $\mathbb{H}\setminus K$ connected is absolutely continuous with respect to $\nu_{\mathbb{H},\kappa}(0,+\infty)$ and that their Radon-Nikodym derivative is explicitly given by
(1) $\frac{\mathrm{d}\nu_{\mathbb{H}\setminus K,\kappa}(0,+\infty)}{\mathrm{d}\nu_{\mathbb{H},\kappa}(0,+\infty)}(\eta)=\mathbf{1}_{\eta\subset\mathbb{H}\setminus K}\varphi^{\prime}(0)^{-h}\exp(\frac{\mathbf{c}}{2}m_{\mathbb{H}}(\eta,K)),$
@@
where $\varphi:\mathbb{H}\setminus K\to\mathbb{H}$ is a conformal equivalence chosen such that $\varphi(0)=0,\varphi(z)\sim z$ as $z\to\infty$, and $h=h_{\kappa}=(6-\kappa)/(2\kappa)$ is a constant that can be related to the weight of a boundary condition changing operator in a CFT [Car05, Pel19, Dub15]. The term $m_{\mathbb{H}}(\eta,K)$ corresponds to the Brownian measure (see [LW04] for a definition thereof) on loops in $\mathbb{H}$ that intersect $K$ and $\eta$. This restriction property is one motivation for seeing the laws $\nu_{D,\kappa}(x,y)$ no longer as probability measures, but rather as measures of total mass [Law09, Law11]
$Z_{D,\kappa}(x,y)=\begin{cases}H_{\partial D}(x,y)^{h}&\text{if }x\neq y\in D\setminus\left\{\infty\right\}\\ \absolutevalue{\varphi^{\prime}(x)}^{h}&\text{if }y=\infty\end{cases}$
where $H_{\partial D}$ denotes the boundary Poisson kernel (when $\partial D$ is sufficiently regular at $x$ and $y$) and $\varphi:D\to\mathbb{H}$ is a conformal equivalence with $\varphi(x)=0$, $\varphi(z)\sim z$ as $z\to\infty$. This total mass is called the partition function of the chordal $\mathrm{SLE}_{\kappa}$ and it is defined up to a constant such that $Z_{\mathbb{H}}(0,+\infty)=1$. The measures $\nu$ then satisfy a form of conformal covariance
$\varphi_{\ast}\nu_{D,\kappa}(x,y)=\absolutevalue{\varphi^{\prime}(x)}^{h}\absolutevalue{\varphi^{\prime}(y)}^{h}\nu_{\varphi(D),\kappa}(\varphi(x),\varphi(y))$
for $x,y\in\partial D$ and $\varphi:D\to\varphi(D)$ a conformal equivalence (smooth at $x$ and $y$), suggesting that these partition functions may be also interpreted as operators in a Conformal field theory [Dub15] and shows they satisfy BPZ equations. This raises a natural question: are the partition functions $Z_{D,\kappa}(x,y)$ positive and finite? If this were the case, the measures could be normalized and the resulting probability measures would be conformally invariant.
@@
Furthermore, Lawler suggested that the restriction property (1) could be used to implicitly but uniquely define $\mathrm{SLE}$ on multiply connected domains $D$ as follows (see Definition 3 for more precision):
@@
Definition 3 (Conformal restriction $\mathrm{SLE}_{\kappa}$ in multiply-connected domains).
@@
For $0\leq\kappa\leq 4$, chordal $\mathrm{SLE}_{\kappa}$ is the family of laws $\left\{\nu_{D,\kappa}(x,y)\right\}_{(D,x,y)}$, where $D\subset\mathbb{C}$ is a finitely-connected domain with smooth boundary at $x$ and $y$, such that
  • the renormalized probability measures are conformally invariant, and
@@
  • the restriction of $\nu_{D,\kappa}(x,y)$ to curves remaining in a simply connected domain $D^{\prime}\subset D$ is absolutely continuous with the $\mathrm{SLE}_{\kappa}$ measure $\nu_{D^{\prime},\kappa}(x,y)$ in $D^{\prime}$, with Radon-Nikodym derivative
    (2) $\frac{\mathrm{d}\nu_{D,\kappa}(x,y)|_{D^{\prime}}}{\mathrm{d}\nu_{D^{\prime},\kappa}(x,y)}(\eta)=\mathbf{1}_{\eta\subset D^{\prime}}\exp(-\frac{\mathbf{c}}{2}m_{D}(\eta,D^{\prime}\setminus D)).$
@@
Because these Radon-Nikodym derivatives are consistent, one can see (for instance via a Carathéodory extension argument) that Equation (2) indeed implicitly defines a ($\sigma$-finite) measure on curves on any finitely connected domain of the complex plane, since the additivity structure of the Brownian loop measure makes the restriction to different domains compatible [Law11]. Moreover, these measures are unique up to a multiplicative constant that one can fix by requiring that $Z_{\mathbb{H}}(0,+\infty)=1$, and satisfy the domain Markov property that is expected of SLE, as well as a form of reversibility. It is further shown in [Law11] that $Z_{D,\kappa}(x,y)$ is positive, and finite either when $\kappa\leq 8/3$ (corresponding to the regime $\mathbf{c}\leq 0$) or when $\kappa\leq 4$ and $D$ is one-connected. The main aim of this paper is to give explicit constructions of the laws satisfying the definition above, and to use those to prove that the partition function is non-trivial.
Before describing our results, let us recall two other ways of constructing $\mathrm{SLE}$ in simply-connected domains, that we will generalize. There is yet a third fruitful way via Liouville quantum gravity and welding of random surfaces [SY23] that we will not treat in this current paper and that is a priori harder to use for the case $\mathrm{SLE}_{4}$ that is the main topic here.
Other constructions of $\mathrm{SLE}$ in simply connected domains
@@
In the case of simply connected domains, $\mathrm{SLE}$ curves can be seen to appear from other conformally invariant objects:
@@
  • The Gaussian Free Field (GFF) corresponds to a centered Gaussian process with covariance operator formally given by Green’s function. Its appearance as a scaling limit for distributions suggests a connection with $\mathrm{SLE}$, which was first discovered in the seminal paper [SS09]. While $\mathrm{SLE}_{\kappa}$ for $\kappa<4$ can be understood as flow lines of the GFF [MS16, She16], it is shown that $\mathrm{SLE}_{4}$ arises as a level-line of the GFF: both as a scaling limit of a level-line of the discrete GFF in [SS09], but also as a purely continuum interface in [SS13]. More precisely, a curve $\eta\sim\nu_{D,4}(x,y)$ in a simply-connected domain $D$ with $x,y\in\partial D$ can be coupled with a Gaussian Free Field $\Phi_{D}$ with mean corresponding to the harmonic extension of the boundary values $-\lambda:=-\sqrt{\pi/8}$ (resp. $+\lambda$) on the counterclockwise arc $(yx)$ (resp. $(xy)$), so that conditionally on $\eta$, $\Phi_{D}$ is equal in law to a sum of two independent GFFs on the left (resp. on the right) of $D\setminus\eta$, with constant mean equal to $-\lambda$ (resp. $+\lambda$). Further works explore variations of this results for other types of SLEs [PW17, HBB10, IK13], and prove the existence of level-lines for GFFs in multiply connected domains [ASW19].
@@
  • The (one-sided) restriction measures and the Conformal Loop Ensembles (CLE) are one-parameter families of measures on hulls (resp. loops) satisfying a form of conformal Markov property. Both can be constructed from Brownian motion: for $D\subset\mathbb{C}$ a bounded domain with $x,y\in\partial D$, the filling (relatively to $(yx)$) of a Poisson point process of Brownian excursions on the boundary arc $(yx)\subset\partial D$ of intensity $\alpha$ yields a one-sided restriction measure of parameter $h=h(\alpha)$ (see Section 3.1 and [LSW03]), whereas the collection of outer boundaries of clusters of loops generated by a Poisson point process of intensity $c$ with respect to the Brownian loop measure corresponds to a $\operatorname{CLE}_{\kappa}$ with $\kappa=\kappa(c)$ (see Section 3.1 and [SW12]). In [WW13], it was shown that for $D\subset\mathbb{C}$ simply connected with $x\neq y\in\partial D$ and $\kappa\in(8/3,4]$, if one samples independently a restriction measure on $(xy)$ of parameter $h_{\kappa}=(6-\kappa)/(2\kappa)$ and a $\operatorname{CLE}$ of parameter $\kappa$, and then proceeds to follow the outer boundary of the restriction sample along with the $\operatorname{CLE}$ loops it intersects, then the curve obtained is distributed like $\mathrm{SLE}_{\kappa}$ in $D$.
@@
In this article, we extend these constructions to the case of multiply connected domains in a way that allows us the study the conformal restriction $\mathrm{SLE}_{\kappa}$ as defined in Definition 3.
In a first part, we provide a construction specific to $\mathrm{SLE}_{4}$ that relies on the connection with the GFF, both when the two target points belong to the same boundary component (non-crossing case) and different boundary components (crossing-case). Loosely speaking, the results for $\mathrm{SLE}_{4}$ can be stated as follows:
@@
    The law of conformal restriction $\mathrm{SLE}_{4}$ in any multiply connected domain can be obtained from a mixture of level-lines of the GFF via appropriate topological conditionings. (See Theorem 913)
@@
    The partition function of conformal restriction $\mathrm{SLE}_{4}$ can be identified, and is finite and positive. (See Corollary 1215)
Somewhat surprisingly, the conformal restriction $\mathrm{SLE}_{4}$ cannot be seen as a level line of a single GFF (see Remark 2.2), except in the crossing-case in the annulus, where it can be interpreted as a level line of a compactified GFF. A bit more precisely, the mixture of laws can be understood as conditioning the final curve on the boundary components it leaves to the left, then sampling a level-line for a GFF whose mean is $-\lambda$ on those boundary components, $+\lambda$ on the others, and conditioning on the level-line to indeed leave those boundary components to the left.
Except for topological discrepancies, the proofs for both cases then follow the same strategy: the law of the level-line $\eta$ of a GFF $\Phi_{D}$ in a domain $D$ with suitable boundary conditions conditioned to remain in a smaller domain $D^{\prime}\subset D$ can be seen as a level-line for a GFF in $D^{\prime}$ with random boundary conditions given by the values $\xi$ of $\Phi_{D}$ on $\partial D^{\prime}$. We first compare this law to $\nu_{D^{\prime},4}(x,y)$ for $\xi$ fixed in Propositions 11 and 17. This builds on the techniques developed in [ALS20b, ALS22] to compare the laws of local sets of GFFs under changes of boundary conditions: Proposition 11 can be seen as a variant of Proposition 3.3 in [ALS22] for more general boundary conditions, but specific to SLEs. In a second step, we compute the average of this Radon-Nikodym derivative over $\xi$, conditionally on $\eta\subset D^{\prime}$. This is a Gaussian integration, and its result consists of a determinant term corresponding to a mass of loops by Proposition 2.2 in [Dub09], and (an exponential of) a Dirichlet energy which we identify as the partition function of the SLE. Equivalently, it can be seen as the inverse extremal distance between parts of the boundary of the domain.
@@
    In any multiply-connected domain and for any $\kappa\in(8/3,4]$ we can construct the conformal restriction $\mathrm{SLE}_{\kappa}$ with the endpoints on different boundary components as a mixture of laws obtained from Brownian excursions and CLE via topological conditioning. (See Theorem 19)
@@
    In the setup just described, one can give an expression for the partition function and show that it is finite. (See Corollary 20)
More precisely, conditionally on the boundary components on the left of the curve, one samples a restriction measure on these boundary components plus the arc $(yx)$ and an independent $\operatorname{CLE}$, and follows the right boundary together with the $\operatorname{CLE}$ loops encountered along the way. To verify that the restriction property (2) is satisfied, it is sufficient to compute the Radon-Nikodym derivative of this law against that in a domain $D\setminus K$, $K$ a hull at positive distance from $(yx)$: the contribution from the $\operatorname{CLE}$ provides the term $\exp(\mathbf{c}\cdot m_{D}(\eta,K)/2)$, whereas the one for the restriction measures yields a normalizing constant that we identify as the partition function of the $\mathrm{SLE}$ (in the simply connected case, this term would be $H_{\partial D}(x,y)^{h}$). We were however not able to extend this construction to the crossing case, which appears to be more delicate - it is not clear how to use the geometric construction via Brownian excursions and CLE to define a crossing $\mathrm{SLE}_{\kappa}$, even in the simplest case of $\mathrm{SLE}_{4}$.
The article is structured as follows. Section 1 reviews standard background on complex analysis tools and $\mathrm{SLE}$. We move the reminders about the GFF and the restriction measures and $\operatorname{CLE}$ to the second and third sections where these tools appear specifically. In Section 2, we construct $\mathrm{SLE}_{4}$ first in the non-crossing case then in the crossing case, from the GFF. Section 3 provides another construction of the $\mathrm{SLE}_{\kappa}$ measure for $\kappa\in(8/3,4]$ via the restriction measures and $\operatorname{CLE}$, but specifically to the non-crossing setting.

### v2

@@ Abstract.
One way to uniquely define Schramm-Loewner Evolution (SLE) in multiply connected domains is to use the conformal restriction property. This gives an implicit definition of a $\sigma$-finite measure on curves; yet it is in general not clear how to construct such measures nor whether the mass of these measures, called the partition function, is finite.
We provide an explicit construction of such SLEs in multiply connected domains when $\kappa=4$ using the Gaussian Free Field (GFF). In particular, we show that there is a mixture of laws of level lines of GFFs that satisfies the restriction property, both when the target points of the curve are on the same or on distinct boundary components. This allows us to give an expression for the partition function of $\mathrm{SLE}_{4}$ on multiply connected domains and shows that the partition function is finite, answering a question raised in [24, 25].
In a second part, we provide a second construction of $\mathrm{SLE}_{\kappa}$ in multiply-connected domains for the whole range $\kappa\in(8/3,4]$: specific, however, to the case of the two target points belonging to the same boundary components. This is inspired by [44] and consists of a mixture of laws on curves obtained by following $\operatorname{CLE}_{\kappa}$ loops and restriction hulls attached to parts of the boundary of the domain. In this case as well, we obtain as a corollary the finiteness of the partition function for this type of $\mathrm{SLE}_{\kappa}$.
@@ Introduction
Schramm-Loewner Evolution $\mathrm{SLE}_{\kappa}$, $\kappa\geq 0$ is a one-dimensional family of conformally invariant probability measures on non-crossing curves in simply connected domains of the complex plane [36] that are either proved or conjectured to describe the scaling limit of interfaces of a variety of planar statistical physics models, e.g. [41, 22, 34, 14, 10, 20, 38]. It also arises in connection with Liouville quantum gravity—see e.g. [16] for a review—and, more relevant to the present work, it is expected to be linked with Conformal Field Theories (CFTs) of central charge $\mathbf{c}=\mathbf{c}(\kappa)=(6-\kappa)(3\kappa-8)/(2\kappa)$, see e.g. [31] for a review and [7, 15] for recent advances in this direction.
In his seminal work, Schramm constructed (chordal) $\mathrm{SLE}_{\kappa}$ in simply connected domains as a solution to Loewner’s equation driven by a multiple of Brownian motion $\sqrt{\kappa}B$ and showed via a concise argument that the resulting law is the unique law on simple curves (modulo increasing time-reparametrisation) that satisfies conformal invariance and a domain Markov property [36]. However, neither the construction via Loewner’s equation nor the characterisation generalise directly to multiply connected domains as the conformal class changes with the growth of the curve and therefore the remaining slit domain cannot be related to the initial one via conformal transformations. In other words, due to the non-triviality of the moduli space for more general domains, identifying a family of laws on curves indexed by the moduli space requires specifying how these laws between different conformal classes compare.
On the other hand, variants of $\mathrm{SLE}_{\kappa}$ have been constructed in doubly connected domains via the annulus Loewner equation using the modulus of the slit domain as the time parametrisation [46, 47, 17, 18], and in higher connectivity using further generalisations of the Loewner equation [6, 5]. In these latter works, one also needs to keep track of the evolution of the moduli as the curve grows in order to obtain a Markov property: the downside is that the natural conditions do not allow to single out a unique law on the curves, yielding instead a whole family of possible curves as $\mathrm{SLE}_{\kappa}$. A different approach that does single out a unique law of $\mathrm{SLE}_{\kappa}$ was suggested by Lawler [24] and is described next.
@@
The (conformal) restriction property describes how the law of SLE in a domain $D$, on the event that it remains in a subdomain $D^{\prime}$, compares to the law of SLE in $D^{\prime}$. More precisely, in the simply connected setting, consider $K\subset\overline{\mathbb{H}}$ compact with $\mathrm{dist}(K,0)>0$ and $\mathbb{H}\setminus K$ (connected and) simply connected. It was shown in [26] that for $\kappa\leq 4$ (the values of $\kappa$ for which the resulting curve is a.s. simple [33]) the law $\nu_{\mathbb{H}\setminus K,\kappa}(0,+\infty)$ of $\mathrm{SLE}_{\kappa}$ from 0 to $\infty$ is absolutely continuous with respect to $\nu_{\mathbb{H},\kappa}(0,+\infty)$ and that their Radon-Nikodym derivative is explicitly given by
[…]
@@
where $\varphi_{K}:\mathbb{H}\setminus K\to\mathbb{H}$ is the conformal equivalence chosen such that $\varphi_{K}(0)=0$ and $\varphi_{K}(z)\sim z$ as $z\to\infty$, and $h=h_{\kappa}=(6-\kappa)/(2\kappa)$ is a constant that can be related to the weight of a boundary condition changing operator in a CFT [9, 31, 13]. The term $m_{\mathbb{H}}(\eta,K)$ corresponds to the Brownian measure (see [23] for a definition thereof) of loops in $\mathbb{H}$ that intersect $K$ and $\eta$. This restriction property is one of the motivations for seeing the laws $\nu_{D,\kappa}(x,y)$ no longer as probability measures, but rather as measures of varying total mass. For $D$ simply connected and $\partial D$ sufficiently regular at $x$ and $y$ this total mass is defined as [24, 25]
[…]
for distinct $x,y\in\partial D$ and $\varphi:D\to\varphi(D)$ a conformal equivalence (that can be extended to the boundary under some assumptions that we detail later on), suggesting that these partition functions may be also interpreted as operators in a Conformal Field Theory [13] satisfying BPZ equations.
@@
Lawler suggested that the restriction property (1) could be used to implicitly but uniquely define $\mathrm{SLE}$ also on multiply connected domains $D$ as follows (see Definition 1.3 for more precision):
@@
Definition 0.1 (Conformal restriction $\mathrm{SLE}_{\kappa}$ in multiply-connected domains).
@@
For $0<\kappa\leq 4$, chordal $\mathrm{SLE}_{\kappa}$ is the family of laws $\left\{\nu_{D,\kappa}(x,y)\right\}_{(D,x,y)}$, where $D\subset\mathbb{C}$ is a finitely-connected domain with smooth boundary at distinct $x,y\in\partial D$, such that
  • conformal covariance is satisfied as in Equation (2)
@@
  • the restriction of $\nu_{D,\kappa}(x,y)$ to curves remaining in a simply connected domain $D^{\prime}\subset D$ is absolutely continuous with respect to the $\mathrm{SLE}_{\kappa}$ measure $\nu_{D^{\prime},\kappa}(x,y)$ in $D^{\prime}$, with Radon-Nikodym derivative
    (3) $\frac{\mathrm{d}\nu_{D,\kappa}(x,y)|_{D^{\prime}}}{\mathrm{d}\nu_{D^{\prime},\kappa}(x,y)}(\eta)=\mathbf{1}_{\eta\subset D^{\prime}}\exp(-\frac{\mathbf{c}}{2}m_{D}(\eta,D\setminus D^{\prime})).$
@@
Because these Radon-Nikodym derivatives are consistent, one can see (for instance via a Carathéodory extension argument) that Equation (3) indeed implicitly defines a ($\sigma$-finite) measure on curves on any finitely connected domain of the complex plane, since the additivity structure of the Brownian loop measure makes the restriction to different domains compatible [25]. Moreover, these measures are unique up to a multiplicative constant fixed by requiring that $Z_{\mathbb{H}}(0,+\infty)=1$, and satisfy the domain Markov property that is expected of SLE, as well as a form of reversibility. This however raises a natural question: are the partition functions, i.e. the total masses $|Z_{D,\kappa}(x,y):=\absolutevalue{\nu_{D,\kappa}(x,y)}$ positive and finite for general multiply-connected domains $D$? If this were the case, the measures could be normalised and the resulting probability measures would be conformally invariant.
Beyond the simply-connected case, it is shown in [25] that $Z_{D,\kappa}(x,y)$ is indeed finite either when $\kappa\leq 8/3$ (corresponding to the regime $\mathbf{c}\leq 0$) or when $\kappa\leq 4$ and $D$ is a conformal annulus. The main aim of this paper is to give explicit constructions of the laws satisfying the definition above, and to use those to prove that the partition function is finite for more general finitely connected domains.
Our results are based on two other ways of constructing $\mathrm{SLE}$ in simply-connected domains, which we will describe next. We will comment on further directions at the end of the introduction.
Two other constructions of $\mathrm{SLE}$ in simply connected domains
@@
In the case of simply connected domains, $\mathrm{SLE}$ curves can be seen to appear from other conformally invariant objects. In particular, the two following connections are instrumental for the present article:
@@
  • As a level line of the Gaussian free field. The Gaussian Free Field (GFF) in a proper domain of the complex plane corresponds to a centred Gaussian process with covariance operator formally given by the Green’s function. Its connection with $\mathrm{SLE}$ was discovered in the seminal paper [34], where it was proved that $\mathrm{SLE}_{4}$ arises as a scaling limit of a certain level-line of the discrete GFF, and in [35] (see also [12]) it was shown how to define the level line purely in the continuum. More precisely, a curve $\eta\sim\nu_{D,4}(x,y)$ in a simply-connected domain $D$ with distinct $x,y\in\partial D$ can be coupled with a Gaussian Free Field $\Phi_{D}$ with mean corresponding to the harmonic extension of the boundary values $\lambda:=-\sqrt{\pi/8}$ (resp. $\lambda$) on the counterclockwise boundary arc $(yx)$ (resp. $(xy)$), so that conditionally on $\eta$, $\Phi_{D}$ is equal in law to a sum of two independent GFFs on the left (resp. on the right) of $D\setminus\eta$, with constant mean equal to $\lambda$ (resp. $\lambda$). Importantly, $\eta$ is then also a measurable function of $\Phi_{D}$. Further works explore variations of this results for other types of SLEs [32, 17, 19, 28], and prove the existence of level-lines for GFFs in multiply connected domains [4].  [^1] In fact, for $\kappa<4$, $\mathrm{SLE}_{\kappa}$ can be also coupled to the GFF as a certain flow line of the GFF [30, 39], but for now these constructions have not been adapted to the multiply-connected setting and it is not clear whether some of our computations would generalise even if they were.
@@
  • Via restriction measures and CLEs. The (one-sided) restriction measures and the Conformal Loop Ensemble (CLE) are one-parameter families of measures on hulls (resp. loops) satisfying a form of conformal Markov property. Both can be constructed from Brownian motion: for $D\subset\mathbb{C}$ a bounded domain with distinct $x,y\in\partial D$, the filling (relatively to $(xy)$) of a Poisson point process of measure proportional (of factor, or intensity $\alpha$) to a measure of Brownian excursions on the boundary arc $(yx)\subset\partial D$ yields a one-sided restriction measure of parameter $h=h(\alpha)$ (see Section 3.1.1 and [26]), whereas the collection of outer boundaries of clusters of loops generated by a Poisson point process of intensity $c$ with respect to the Brownian loop measure corresponds to a $\operatorname{CLE}_{\kappa}$ with $\kappa=\kappa(c)$ (see Section 3.1.2 and [37]). In [44], it was shown that for $D\subset\mathbb{C}$ a simply connected domain with distinct $x,y\in\partial D$ and $\kappa\in(8/3,4]$, if one samples independently a restriction measure on $(yx)$ of parameter $h_{\kappa}=(6-\kappa)/(2\kappa)$ and a $\operatorname{CLE}$ of parameter $\kappa$, and then proceeds to follow the outer boundary of the restriction sample along with the $\operatorname{CLE}$ loops it intersects, then the curve obtained is distributed like $\mathrm{SLE}_{\kappa}$ in $D$.
@@
In this article, we extend these connections to the case of multiply connected domains in a way that provides an explicit construction of the $\mathrm{SLE}_{\kappa}$ described by Definition 0.1. We sometimes refer to this measure as conformal restriction SLE to emphasise that a priori different properties might be used to define $\mathrm{SLE}_{\kappa}$ in multiply connected domains and they might give rise to different measures.
In the first part, we provide a construction specific to $\mathrm{SLE}_{4}$ that relies on the connection with the GFF, both when the two target points (i.e. start and end-point of SLE) belong to the same boundary component (non-crossing case) or different boundary components (crossing-case). Informally, the results for $\mathrm{SLE}_{4}$ can be stated as follows:
@@
    The law of conformal restriction $\mathrm{SLE}_{4}$ in any multiply connected domain can be obtained from a mixture of level-lines of the GFF via appropriate topological conditionings. (See Theorem 2.6, 2.12)
@@
    The partition function of conformal restriction $\mathrm{SLE}_{4}$ can be identified, and is finite and positive. (See Corollary 2.7, 2.13)
Somewhat surprisingly, except in the crossing-case in the annulus where it can be interpreted as a level line of a compactified GFF, the conformal restriction $\mathrm{SLE}_{4}$ cannot in general be seen as a level line of a single GFF, but is rather given as a mixture of conditioned level lines. For instance in the non-crossing case, a simple curve connecting distinct $x,y\in\partial D$ lying on the same boundary component partitions the other boundary components into $\Upsilon_{l}\cup\Upsilon_{r}$ where $\Upsilon_{l}$ is the set of those on its left and $\Upsilon_{r}$ the set of those on its right. The mixture of laws yielding $\mathrm{SLE}_{4}$ can then be understood as a mixture of certain conditional laws: for each such partition, we sample a level-line for a GFF whose boundary value is $\lambda$ on $\Upsilon_{l}$, $\lambda$ on $\Upsilon_{r}$, and finally condition this level-line to indeed partition the boundary components into exactly $\Upsilon_{l}\cup\Upsilon_{r}$. Such a mixture cannot be obtained by just choosing appropriate boundary conditions upfront, see Remark 2.11.
Except for topological discrepancies, the proofs for both cases then follow the same strategy: the law of the level-line $\eta$ of a GFF $\Phi_{D}$ in a domain $D$ with suitable boundary conditions conditioned to remain in a smaller domain $D^{\prime}\subset D$ can be seen as a level-line for a GFF in $D^{\prime}$ with random boundary conditions given by the values $\xi$ of $\Phi_{D}$ on $\partial D^{\prime}$. We first compare this law to $\nu_{D^{\prime},4}(x,y)$ for $\xi$ fixed in Propositions 2.9 and 2.14. This builds on the techniques developed in [2, 3] to compare the laws of local sets of GFFs under changes of boundary conditions: Proposition 2.9 can be seen as a variant of Proposition 3.3 in [3] for more general boundary conditions, but specific to SLEs. In a second step, we compute the average of this Radon-Nikodym derivative over $\xi$, conditionally on $\eta\subset D^{\prime}$. This is a Gaussian integration, and its result consists of a determinant term corresponding to a mass of loops by Proposition 2.2 in [12], and (an exponential of) a Dirichlet energy which we identify as the partition function of the SLE. Equivalently, it can be seen as the inverse extremal distance between parts of the boundary of the domain.
@@
    In any multiply-connected domain and for any $\kappa\in(8/3,4]$ we can construct the conformal restriction $\mathrm{SLE}_{\kappa}$ with the endpoints on the same boundary components as a mixture of laws obtained from Brownian excursions and CLE via topological conditioning. (See Theorem 3.2)
@@
    In the setup just described, one can give an expression for the partition function and show that it is finite. (See Corollary 3.3)
More precisely, conditionally on the boundary components on the left of the curve, one samples a Poisson point process of Brownian excursions on these boundary components plus the arc $(yx)$ and an independent $\operatorname{CLE}$, and follows the right boundary together with the $\operatorname{CLE}$ loops encountered along the way. To verify that the restriction property (3) is satisfied, it is sufficient to compute the Radon-Nikodym derivative of this law against that in a domain $D\setminus K$, $K$ a hull at positive distance from $(yx)$: the contribution from the $\operatorname{CLE}$ provides the term $\exp(\mathbf{c}\cdot m_{D}(\eta,K)/2)$, whereas the one for the Brownian excursions yields a normalising constant that we identify as the partition function of the $\mathrm{SLE}$ (in the simply connected case, this term would be $H_{\partial D}(x,y)^{h}$). We were however not able to extend this construction to the crossing case, which appears to be more delicate - it is not clear how to use the geometric construction via Brownian excursions and CLE to define a crossing $\mathrm{SLE}_{\kappa}$, even in the simplest case of $\mathrm{SLE}_{4}$.
Further constructions and connections
We now comment on related work concerning SLEs in non-simply connected domains and their potential connections to lattice models.
First, parallel to Lawler’s work [25] showing finiteness of the partition function on annuli, Zhan [46, 47] introduces a family of processes $\mathrm{SLE}(\kappa,\Lambda_{\kappa;\left\langle s\right\rangle})$ as the unique law on curves in annuli satisfying the domain Markov property and reversibility, and shows [48] that conformal restriction is also satisfied under a setting slightly different from the above. In fact, it is argued in [25] that both constructions coincide when the laws are normalised. [19] also introduce a variety of SLE curves in the annulus as level-lines of GFFs with various boundary conditions and describe their driving function relative to the annulus Loewner equation, but the conformal restriction properties are not discussed.
Second, the restriction property also appears in the study of discrete models conjectured to converge to $\mathrm{SLE}$ in general domains. An example of such a model is the $\lambda$-Self Avoiding Walks (SAW), introduced in [21]. Already in the discrete, it satisfies a certain random-walk related restriction property and this property should give rise to the the restriction property described above in the scaling limit. See [25] for conjectures related to the $\lambda$-SAW and $\mathrm{SLE}_{\kappa}$, which should also hold in the non-simply-connected setting. The case of the more standard lattice models is less clear—for example it not obvious which Ising interface, or with which boundary conditions, would converge either to the crossing or non-crossing annulus $\mathrm{SLE}_{3}$ defined above. We find it is an interesting question to figure it out.
Finally, there is yet a third fruitful way via Liouville quantum gravity and welding of random surfaces [42] that we will not treat in this current paper (and that is a priori harder to use for the case $\mathrm{SLE}_{4}$ that is the main topic here).
Structure of the article
The article is structured as follows. Section 1 reviews standard background on complex analysis tools and $\mathrm{SLE}$. We move the reminders about the GFF, resp. the restriction measures and $\operatorname{CLE}$ to the second and third sections where these tools appear specifically. In Section 2, we construct $\mathrm{SLE}_{4}$ first in the non-crossing case then in the crossing case, from the GFF. Section 3 provides another construction of the $\mathrm{SLE}_{\kappa}$ measure for $\kappa\in(8/3,4]$ via the Brownian excursions and $\operatorname{CLE}$, but specific to the non-crossing setting.